# SVarM: Linear Support Varifold Machines for Classification and Regression on Geometric Data

### v1

@@ Abstract
Despite progress in the rapidly developing field of geometric deep learning, performing statistical analysis on geometric data—where each observation is a shape such as a curve, graph, or surface—remains challenging due to the non-Euclidean nature of shape spaces, which are defined as equivalence classes under invariance groups. Building machine learning frameworks that incorporate such invariances, notably to shape parametrization, is often crucial to ensure generalizability of the trained models to new observations. This work proposes SVarM to exploit varifold representations of shapes as measures and their duality with test functions $h:\mathbb{R}^{n}\times S^{n-1}\rightarrow\mathbb{R}$. This method provides a general framework akin to linear support vector machines but operating instead over the infinite-dimensional space of varifolds. We develop classification and regression models on shape datasets by introducing a neural network-based representation of the trainable test function $h$. This approach demonstrates strong performance and robustness across various shape graph and surface datasets, achieving results comparable to state-of-the-art methods while significantly reducing the number of trainable parameters.
@@ 1 Introduction
Shape analysis, or geometric data science, is a field dedicated to building statistical and machine learning methods able to retrieve and analyze the morphological variability in geometric structures. This is a particularly central problem in applications such as computer vision or biomedical imaging where observations often come as segmented curves, surfaces, densities or other types of complex geometric data. Various approaches have been proposed, including Riemannian and elastic shape space models [21, 36, 27, 51], topology based methods [20, 13, 18], and metric/functional matching frameworks [9, 35, 37]. With the recent advances in acquisition protocols and the explosion in the size and resolution of datasets that followed, one can expect deep learning to eventually provide tools that better scale up to modern data dimensionality. Despite this, the very nature of geometric data poses unique challenges when it comes to designing robust neural network models to tackle shape analysis tasks, which is the object of active research within the relatively new area known as geometric deep learning, see the surveys [10, 12, 19] and references therein for the current state-of-the-art.
@@
In this work, we introduce a shape learning framework based on a different construction than aforementioned references. The main idea is to leverage the concept of varifolds from geometric measure theory [2, 1] along with the natural representation of a shape $q\in\mathcal{S}$—such as a curve or a surface in $\mathbb{R}^{n}$—as its corresponding varifold $\mu_{q}\in\mathcal{V}$. The mapping $q\mapsto\mu_{q}$ is invariant under reparameterization, thereby providing a (nonlinear) embedding of the shape space $\mathcal{S}$ into $\mathcal{V}$, the space of measures on $\mathbb{R}^{n}\times S^{n-1}$. This enables us to instead build the learning framework on the space of varifolds by exploiting the duality of these measures with test functions. We here simply propose to rely on linear (or affine) forms of $\mathcal{V}$ of the form $\langle\mu,h\rangle$, where $h\in C_{0}(\mathbb{R}^{n}\times S^{n-1})$ is a trainable test function. Due to its analogy with support vector machines (but here considered in the infinite-dimensional varifold space), we coin our model Support Varifold Machines (SVarM) and derive formulations of the shape regression and classification problems based on this setting. Importantly, this approach allows for a principled handling of complex shapes without requiring prior regularization or registration. Additionally, by focusing on approximating continuous functions in a low-dimensional space, the model remains lightweight, requiring significantly fewer trainable parameters than traditional models operating on raw geometric data as inputs.
@@
The presentation and contributions of the paper are organized into three main sections. Section 2 introduces the mathematical settings of shape spaces and of varifolds. We then proceed by deriving a few theoretical results related to the approximation properties of SVarM for regression and classification as well as its stability with respect to geometric perturbations. Section 3 details our proposed neural network architecture, which approximates the trainable test function $h$ of the model. Finally, Section 4 demonstrates the practical effectiveness of the approach on various regression and classification problems for real surface and shape graph datasets.
@@ 2.1 Shape Space
The concept of ”shape” encompasses a wide variety of geometric structures (see e.g. [51, 5, 44] for general overviews on shape analysis). Although the framework we introduce can in principle apply to any type of shapes that can be viewed as measures or varifolds, our focus here will be on open/closed curves embedded in the plane or 3D space, surfaces in $\mathbb{R}^{3}$, and by extension structures known as shape graphs [46, 3]. The challenging aspect of dealing with such objects is that there is generally not a simple nor canonical way to represent them in a common Euclidean or even functional space. Indeed, a curve or surface is usually defined by a parametrization function $q:M\rightarrow\mathbb{R}^{n}$, where $n=2$ or $3$ and $M$ is the parameter manifold (for instance an interval or $S^{1}$ for open and closed curves, a disk or $S^{2}$ for open/closed $\ldots$). In this paper, we will consider rectifiable shapes where $q$ is assumed to be a bi-Lipschitz regular embedding from $M$ to $\mathbb{R}^{n}$, denoted by $q\in\mathcal{I}\doteq\operatorname{Emb}_{\text{Lip}}(M,\mathbb{R}^{n})$. However, $q$ itself does not uniquely identify the shape $q(M)\subseteq\mathbb{R}^{n}$. The action of any reparametrization $q\circ\gamma$, where $\gamma\in\mathcal{D}=\operatorname{Diff}_{\text{Lip}}(M)$ is a bi-Lipschitz homeomorphism of $M$, leaves the shape $q(M)$ invariant. In other words, the actual shape associated to $q$ is in essence the equivalence class $[q]=\{q\circ\gamma:\ \gamma\in\operatorname{Diff}_{\text{Lip}}(M)\}$, and the shape space $\mathcal{S}$ can be thus identified with the quotient $\mathcal{S}=\mathcal{I}/\mathcal{D}$, see Figure 1 for an illustration. Note that quotienting out other shape invariances, such as to rigid motions of $\mathbb{R}^{n}$, can also be relevant in certain situations, but this work focuses on creating a parametrization-invariant model, as rigid alignment can usually be applied in a separate step. In fact, Section 4.1 and Section 4.2 illustrates how SVarM can itself be used for that very task.
@@ 2.2 Varifolds
We shall consider the following particular and simpler version of the notion of varifolds introduced in [1], similar to the framework of [28], which is specifically tailored to shapes of dimension or co-dimension $1$ such as curves and surfaces that are the focus of this paper:
@@
[…]
Now, going back to shape spaces, any Lipschitz embedded curve or surface $q:M\rightarrow\mathbb{R}^{n}$ (where $M$ is of dimension $1$ or $n-1$) can be canonically mapped to a varifold $\mu_{q}\in\mathcal{V}$ that is defined as the pushforward measure $\mu_{q}:=(q,v_{q})_{*}\operatorname{vol}_{q}$, where $v_{q}$ denotes the unit tangent/normal vector field and $\operatorname{vol}_{q}$ is the induced length/area measure on $M$. More explicitly, for any $h\in C_{0}(\mathbb{R}^{n}\times S^{n-1})$,
@@
For any $q\in\mathcal{I}$ and $\gamma\in\mathcal{D}$, it holds that $\mu_{q\circ\gamma}=\mu_{q}$. Consequently, one obtains a well-defined map $[q]\in\mathcal{S}\mapsto\mu_{q}\in\mathcal{V}$ which, in addition, is injective.
@@
[…]
where $\mathcal{H}^{d}$ denotes the $d$-dimensional Hausdorff measure of $\mathbb{R}^{n}$, and $v_{S}(x)$ is the tangent/normal vector to $S$ at $x$ which is defined for $\mathcal{H}^{d}$-a.a. $x\in S$. This shows that $\langle\mu_{q},h\rangle$ only depends on $S=q(M)$ and not the parametrization $q$ itself.
Next we need to prove the injectivity of the varifold map on $\mathcal{S}$. Let $[q]$, $[q^{\prime}]$ be two distinct shapes of $\mathcal{S}$. Then $S=q(M)$ and $S^{\prime}=q^{\prime}(M)$ are two distinct rectifiable subsets of dimension $d$ of $\mathbb{R}^{n}$. Without loss of generality, we may assume that $\mathcal{H}^{d}(S\setminus S^{\prime})>0$. By continuity, one can deduce that there exist $x_{0}\in S$, $r>0$ and $\epsilon>0$ such that $\text{dist}((B(x_{0},r),S^{\prime})>\epsilon>0$. Then one can construct a nonnegative test function $h\in C_{0}(\mathbb{R}^{n}\times S^{n-1})$ such that $h(x,v)=1$ for any $x\in B(x_{0},r)$ and $h(x,v)=0$ for $\|x-x_{0}\|\geq r+\epsilon$. From (2), we obtain that:
[…]
@@ 2.3.1 Existence of Regression Functions and Separators
In particular, for a varifold $\mu_{q}$ associated to $q\in\mathcal{I}$, one can see from (2) that $\|\mu_{q}\|_{TV}$ is the total length/area of the shape $q(M)$. The TV norm induces a rather strong topology on varifolds: for instance, the TV norm between two Dirac masses $\delta_{(x_{0},v_{0})}$ and $\delta_{(x_{1},v_{1})}$ is $2$ whenever $(x_{0},v_{0})\neq(x_{1},v_{1})$ and 00 when the position and direction coincide. It is often more natural to consider the weak-* topology on $\mathcal{M}(\mathbb{R}^{n}\times S^{n-1})$. By definition, a sequence $(\mu_{i})$ weak-* converges to $\mu$, denoted $\mu_{i}\buildrel\ast\over{\rightharpoonup}\mu$, when $\langle\mu_{i},h\rangle\xrightarrow[i\rightarrow\infty]{}\langle\mu,h\rangle$ for all $h\in C_{0}(\mathbb{R}^{n}\times S^{n-1})$. Note that the weak-* convergence can be also locally metrized by a variety of usual metrics on $\mathcal{M}(\mathbb{R}^{n}\times S^{n-1})$ such as the bounded Lipschitz norm or the Wasserstein metrics of optimal transport (for normalized measures), c.f. [48] Chap. 6.
@@
Without loss of generality, let us assume that $C_{1},C_{2}\subseteq\mathcal{S}$ are two finite sets in $\mathcal{V}$ with $C_{1}=\{\mu_{1},\ldots,\mu_{p}\}$ and $C_{2}=\{\nu_{1},\ldots,\nu_{q}\}$, such that for each $\nu\in C_{2}$, $\text{supp}(\nu)\not\subseteq\bigcup_{\mu\in C_{1}}\text{supp}(\mu)$. Let us denote by $\text{conv}(C_{1})$ and $\text{conv}(C_{2})$ the convex hulls of the two sets $C_{1}$ and $C_{2}$ in $\mathcal{V}$. By contradiction, if $\text{conv}(C_{1})\cap\text{conv}(C_{2})\neq\emptyset$, there exist $(\alpha_{1},\ldots,\alpha_{p})$ and $(\beta_{1},\ldots,\beta_{q})$ in the probability simplexes of +$\mathbb{R}^{p}_{+}$ and $\mathbb{R}_{+}^{q}$ respectively such that:
@@
Since not all $\beta_{j}$’s can be 00, we deduce that there exist $j$ such that $\operatorname{supp}(\nu_{j})\subseteq\bigcup_{\mu\in C_{1}}\operatorname{supp}(\mu)$ which contradicts the assumption. Therefore, $\text{conv}(C_{1})$ and $\text{conv}(C_{2})$ are two disjoint convex sets in $\mathcal{V}$. They are also weak-* compact as the convex hulls of finite sets; indeed $\text{conv}(C_{1})$ is the image of the compact probability simplex by the continuous map $(\alpha_{1},\ldots,\alpha_{p})\mapsto\sum_{i=1}^{p}\alpha_{i}\mu_{i}$, and similarly for $\text{conv}(C_{2})$. Then, by Theorem 2, we deduce that $\text{conv}(C_{1})$ and $\text{conv}(C_{2})$ and thus $C_{1}$ and $C_{2}$ can be separated by a test function $h\in C_{0}(\mathbb{R}^{n}\times S^{n-1})$. ∎
@@ 2.3.2 Stability with Respect to Shape Deformations and Perturbations
Another key theoretical question regarding the SVarM approach is the stability of the model under perturbations of the input data. We show below that this type of result can hold under the assumption that $h$ is Lipchitz continuous. As described in Section 3, we will approximate $h$ using a multilayer perceptron (MLP) architecture. For fixed weights (i.e., after training), an MLP is a Lipchitz function, and one can efficiently compute an upper bound on its Lipchitz constant [49]. Let us first examine stability for general varifolds in $\mathcal{V}$:
@@
Next we define the function $\overline{h}:=\frac{h}{K}$. Since the Lipchitz constant of $h$ is $K$, it follows that $\overline{h}$ is $1-$Lipchitz. Thus we can write
@@
Let $h\in C_{0}(\mathbb{R}^{n}\times S^{n-1})$ such that $h$ is Lipschitz with constant $K$ and $q,q^{\prime}$ be bi-Lipschitz embeddings from a $d$-dimensional manifold $M$ into $\mathbb{R}^{n}$ (with either $d=1$ or $d=n-1$) such that $\|q-q^{\prime}\|_{W^{1,\infty}}\leq C$. Then
@@
3 Model Architecture

### v2

@@ Abstract
Despite progress in the rapidly developing field of geometric deep learning, performing statistical analysis on geometric data—where each observation is a shape such as a curve, graph, or surface—remains challenging due to the non-Euclidean nature of shape spaces, which are defined as equivalence classes under invariance groups. Building machine learning frameworks that incorporate such invariances, notably to shape parametrization, is often crucial to ensure generalizability of the trained models to new observations. This work proposes SVarM to exploit varifold representations of shapes as measures and their duality with test functions $h:\mathbb{R}^{n}\times S^{n-1}\rightarrow\mathbb{R}$. This method provides a general framework akin to linear support vector machines but operating instead over the infinite-dimensional space of varifolds. We develop classification and regression models on shape datasets by introducing a neural network-based representation of the trainable test function $h$. This approach demonstrates strong performance and robustness across various shape graph and surface datasets, achieving results comparable to state-of-the-art methods while significantly reducing the number of trainable parameters. Keywords: geometric deep learning, linear regression, classification, varifolds.
@@ 1 Introduction
Shape analysis, or geometric data science, is a field dedicated to building statistical and machine learning methods able to retrieve and analyze the morphological variability in geometric structures. This is a particularly central problem in applications such as computer vision or biomedical imaging where observations often come as segmented curves, surfaces, densities or other types of complex geometric data. Various approaches have been proposed, including Riemannian and elastic shape space models [23, 41, 31, 56], topology based methods [22, 15, 20], and metric/functional matching frameworks [11, 40, 42]. These different methods have proved quite successful in tackling problems such as pairwise comparison, regression, classification or clustering for datasets of shapes. However, with the constant advances in acquisition protocols and the explosion in the size and resolution of datasets that followed, many such methods do not always scale well to recent applications that may involve databases with up to tens of thousands of subjects, each made of potentially hundreds of thousands of vertices. In view of the rapid development of new machine learning paradigms, in particular neural network models, and their impressive achievements in image processing and analysis tasks, one can reasonably expect similar tools to be able to address those challenges on geometric data. Yet, the very particular and intricate nature of shape spaces poses unique challenges when it comes to designing robust neural network models for shape analysis tasks. The importance of this problem within machine learning was recently recognized through the emergence of the research area known geometric deep learning, see the surveys [12, 14, 21] and references therein for the current state-of-the-art.
@@
In contrast with many works in geometric deep learning that focus on developing neural network architectures (in particular convolutional models) defined directly on e.g. surface meshes [34, 32, 24], the approach we propose in this paper instead relies on the embedding of shapes into a certain measure space. Specifically, our learning framework leverages the concept of varifolds from geometric measure theory [2, 1] along with the natural representation of a shape $q\in\mathcal{S}$—such as a curve or a surface in $\mathbb{R}^{n}$—as its corresponding varifold $\mu_{q}\in\mathcal{V}$. The mapping $q\mapsto\mu_{q}$ is invariant under reparameterization, thereby providing a (nonlinear) embedding of the shape space $\mathcal{S}$ into $\mathcal{V}$, the space of measures on $\mathbb{R}^{n}\times S^{n-1}$. This enables us to instead build the learning framework on the space of varifolds by exploiting the duality of these measures with test functions. We here simply propose to rely on affine functions of $\mathcal{V}$ of the form $\langle\mu,h\rangle+\beta$, where $h\in C_{0}(\mathbb{R}^{n}\times S^{n-1})$ is a trainable test function and $\beta\in\mathbb{R}$ a trainable bias. Due to its analogy with support vector machines (but here considered in the infinite-dimensional varifold space), we coin our model Support Varifold Machines (SVarM) and derive formulations of the shape regression and classification problems based on this setting. This overall idea also shares certain similarities with recent works such as [8] which introduces a regression model for function data based on the square root velocity function (SRVF) representation. Our use of varifolds provides, arguably, a more convenient setting when dealing with geometric instead of functional data. Importantly, our approach allows for a principled handling of complex shapes without requiring prior regularization or registration. Additionally, by focusing on approximating continuous functions in a low-dimensional space, the model remains lightweight, requiring significantly fewer trainable parameters than traditional models operating on raw geometric data as inputs.
@@
The presentation and contributions of the paper are organized into three main sections. Section 2 introduces the mathematical settings of shape spaces and of varifolds. We then proceed by deriving a few theoretical results related to the approximation properties of SVarM for regression and classification as well as its stability with respect to geometric perturbations. Section 3.2 details our proposed neural network architecture, which approximates the trainable test function $h$ of the model. Finally, Section 4 demonstrates the practical effectiveness of the approach on various regression and classification problems for real surface and shape graph datasets.
@@ 2.1 Shape Space
[…]
@@ 2.2 Varifolds
To be exact, the above definition corresponds to dimension/codimension 1 oriented varifolds of $\mathbb{R}^{n}$; a more general construction can be obtained by replacing $S^{n-1}$ with Grassmannians, c.f. [1, 17]. We will then denote $\text{supp}(\mu)\subseteq\mathbb{R}^{n}\times S^{n-1}$ the support of a varifold $\mu\in\mathcal{V}$. Furthermore, by the usual Riesz-Markov-Kakutani representation theorem, any varifold $\mu$ can be viewed as an element of the dual to $C_{0}(\mathbb{R}^{n}\times S^{n-1})$, the space of continuous test functions on $\mathbb{R}^{n}\times S^{n-1}$ vanishing at infinity and we shall use the notation $\langle\mu,h\rangle=\int_{\mathbb{R}^{n}\times S^{n-1}}h(x,v)d\mu(x,v)$ for the duality bracket between $\mathcal{V}$ and $C_{0}(\mathbb{R}^{n}\times S^{n-1})$. In particular, a Dirac varifold $\delta_{(x_{0},v_{0})}$ for some $(x_{0},v_{0})\in\mathbb{R}^{n}\times S^{n-1}$ is such that $\langle\delta_{(x_{0},v_{0})},h\rangle=h(x_{0},v_{0})$.
Now, going back to shape spaces, any Lipschitz embedded curve or surface $q:M\rightarrow\mathbb{R}^{n}$ (where $M$ is of dimension $1$ or n1n-1) can be canonically mapped to a varifold $\mu_{q}\in\mathcal{V}$ that is defined as the pushforward measure $\mu_{q}:=(q,v_{q})_{*}\operatorname{vol}_{q}$, where $v_{q}$ denotes the unit tangent/normal vector field and $\operatorname{vol}_{q}$ is the induced length/area measure on $M$. More explicitly, for any $h\in C_{0}(\mathbb{R}^{n}\times S^{n-1})$,
@@
For any $q\in\mathcal{I}$ and $\gamma\in\mathcal{D}$, it holds that $\mu_{q\circ\gamma}=\mu_{q}$. Consequently, one obtains a well-defined map $q]\in\mathcal{S}\mapsto\mu_{q}\in\mathcal{V}$ which, in addition, is injective.
@@
$\langle\mu_{q},h\rangle=\int_{M}h(q(m),v_{q}(m))d\,\text{vol}_{q}(m)=\int_{S}h(x,v_{S}(x))d\mathcal{H}^{d}(x)$ (2)
where $\mathcal{H}^{d}$ denotes the $d$-dimensional Hausdorff measure of $\mathbb{R}^{n}$, and $v_{S}(x)$ is the tangent/normal vector to $S$ at $x$ which is defined for $\mathcal{H}^{d}$-$a.$ $x\in S$. This shows that $\langle\mu_{q},h\rangle$ only depends on $S=q(M)$ and not the parametrization $q$ itself.
Next we need to prove the injectivity of the varifold map on $\mathcal{S}$. Let $[q]$, $q^{\prime}]$ be two distinct shapes of $\mathcal{S}$. Then $S=q(M)$ and $S^{\prime}=q^{\prime}(M)$ are two distinct rectifiable subsets of dimension $d$ of $\mathbb{R}^{n}$. Without loss of generality, we may assume that $\mathcal{H}^{d}(S\setminus S^{\prime})>0$. By continuity, one can deduce that there exist $x_{0}\in S$, $r>0$ and $\epsilon>0$ such that $\text{dist}((B(x_{0},r),S^{\prime})>\epsilon>0$. Then one can construct a nonnegative test function $h\in C_{0}(\mathbb{R}^{n}\times S^{n-1})$ such that $h(x,v)=1$ for any $x\in B(x_{0},r)$ and $h(x,v)=0$ for $\|x-x_{0}\|\geq r+\epsilon$. From (2), we obtain that:
$\langle\mu_{q},h\rangle=\int_{S}h(x,v_{S}(x))d\mathcal{H}^{d}(x)\geq\int_{B(x_{0},r)\cap S}h(x,v_{S}(x))d\mathcal{H}^{d}(x)=\mathcal{H}^{d}(B(x_{0},r)\cap S)>0$
@@ 2.3.1 Existence of Regression Functions and Separators
In particular, for a varifold $\mu_{q}$ associated to $q\in\mathcal{I}$, one can see from (2) that $\|\mu_{q}\|_{TV}$ is the total length/area of the shape $q(M)$. The TV norm induces a rather strong topology on varifolds: for instance, the TV norm between two Dirac masses $\delta_{(x_{0},v_{0})}$ and $\delta_{(x_{1},v_{1})}$ is $2$ whenever $(x_{0},v_{0})\neq(x_{1},v_{1})$ and 0 when the position and direction coincide. It is often more natural to consider the weak-* topology on $\mathcal{M}(\mathbb{R}^{n}\times S^{n-1})$. By definition, a sequence $(\mu_{i})$ weak-* converges to $\mu$, denoted $\mu_{i}\buildrel\ast\over{\rightharpoonup}\mu$, when $\langle\mu_{i},h\rangle\xrightarrow[i\rightarrow\infty]{}\langle\mu,h\rangle$ for all $h\in C_{0}(\mathbb{R}^{n}\times S^{n-1})$. Note that the weak-* convergence can be also locally metrized by a variety of usual metrics on $\mathcal{M}(\mathbb{R}^{n}\times S^{n-1})$ such as the bounded Lipschitz norm or the Wasserstein metrics of optimal transport (for normalized measures), c.f. [53] Chap. 6.
@@
Without loss of generality, let us assume that $C_{1},C_{2}\subseteq\mathcal{S}$ are two finite sets in $\mathcal{V}$ with $C_{1}=\{\mu_{1},\ldots,\mu_{p}\}$ and $C_{2}=\{\nu_{1},\ldots,\nu_{q}\}$, such that for each $\nu\in C_{2}$, $\text{supp}(\nu)\not\subseteq\bigcup_{\mu\in C_{1}}\text{supp}(\mu)$. Let us denote by $\text{conv}(C_{1})$ and $\text{conv}(C_{2})$ the convex hulls of the two sets $C_{1}$ and $C_{2}$ in $\mathcal{V}$. By contradiction, if $\text{conv}(C_{1})\cap\text{conv}(C_{2})\neq\emptyset$, there exist $(\alpha_{1},\ldots,\alpha_{p})$ and $(\beta_{1},\ldots,\beta_{q})$ in the probability simplexes of $\mathbb{R}^{p}_{+}$ and $\mathbb{R}_{+}^{q}$ respectively such that:
@@
Since not all $\beta_{j}$’s can be 0, we deduce that there exist $j$ such that $\operatorname{supp}(\nu_{j})\subseteq\bigcup_{\mu\in C_{1}}\operatorname{supp}(\mu)$ which contradicts the assumption. Therefore, $\text{conv}(C_{1})$ and $\text{conv}(C_{2})$ are two disjoint convex sets in $\mathcal{V}$. They are also weak-* compact as the convex hulls of finite sets; indeed $\text{conv}(C_{1})$ is the image of the compact probability simplex by the continuous map $(\alpha_{1},\ldots,\alpha_{p})\mapsto\sum_{i=1}^{p}\alpha_{i}\mu_{i}$, and similarly for $\text{conv}(C_{2})$. Then, by Theorem 2, we deduce that $\text{conv}(C_{1})$ and $\text{conv}(C_{2})$ and thus $C_{1}$ and $C_{2}$ can be separated by a test function $h\in C_{0}(\mathbb{R}^{n}\times S^{n-1})$. ∎
@@ 2.3.2 Stability with Respect to Shape Deformations and Perturbations
Another key theoretical question regarding the SVarM approach is the stability of the model under perturbations of the input data. We show below that this type of result can hold under the assumption that $h$ is Lipchitz continuous. As described in Section 3.2, we will approximate $h$ using a multilayer perceptron (MLP) architecture. For fixed weights (i.e., after training), an MLP is a Lipchitz function, and one can efficiently compute an upper bound on its Lipchitz constant [54]. Let us first examine stability for general varifolds in $\mathcal{V}$:
@@
Next we define the function $\overline{h}:=\frac{h}{K}$. Since the Lipchitz constant of $h$ is $K$, it follows that $\overline{h}$ is 11-Lipchitz. Thus we can write
@@
Let $h\in C_{0}(\mathbb{R}^{n}\times S^{n-1})$ such that $h$ is Lipschitz with constant $K$ and $q,q^{\prime}$ be bi-Lipschitz embeddings from a $d$-dimensional manifold $M$ into $\mathbb{R}^{n}$ (with either $d=1$ or d=n1d=n-1) such that $\|q-q^{\prime}\|_{W^{1,\infty}}\leq C$. Then
@@
3 SVarM regression and classification models
The core of the SVarM approach for regression or classification is to approximate a map $h:\mathbb{R}^{n}\times S^{n-1}\to\mathbb{R}$ (for $n=2$ or $3$) which, together with the bias $\beta$, represents any affine map on $\mathcal{V}$ as we have previously shown in Proposition 1. In the absence of any clear finite-dimensional family of functions to represent $h$, we focus here on non-parametric models; we first examine the situation in which $h$ is taken in some reproducing kernel Hilbert space of functions on $\mathbb{R}^{n}\times S^{n-1}$ leading to a formulation akin to kernel regression/classification, before relying instead on the representation of $h$ via a neural network model which we argue is a more effective approach when dealing with large datasets. From now on, we shall assume that a training set $\{\mu^{(k)},y^{(k)}\}_{k=1,\ldots,R}$ is given, where $\{\mu^{(1)},...,\mu^{(R)}\}$ is a set of varifolds (computed from the corresponding input shapes) and $\{y^{(1)},...,y^{(R)}\}$ are the corresponding labels (either real numbers in regression problems or integers giving class membership for classification).
3.1 Reproducing kernel Hilbert space approach
Reproducing kernel Hilbert spaces (RKHS) in the context of varifolds have been considered in several previous works [17, 33, 29, 43], primarily as a way to build explicit metrics on the space of curves and surfaces. One can rely on a similar construction in order to specify the space of test functions $h$, in which case the problem is related to kernel ridge regression with a theoretically explicit form for the solution. Before discussing this approach, we begin with a brief recap of the main background material on RKHS, referring the reader to e.g. [3] or [56] Chap. 8 for more detailed presentations.
Let us start by introducing a positive definite kernel function $K$ on $\mathbb{R}^{n}\times S^{n-1}$, i.e. a function $K:(\mathbb{R}^{n}\times S^{n-1})\times(\mathbb{R}^{n}\times S^{n-1})\rightarrow\mathbb{R}$ wich is symmetric and satisfies the property that for any $(x_{i},v_{i})_{i=1,\ldots,m}\in(\mathbb{R}^{n}\times S^{n-1})^{m}$ and $(w_{i})\in\mathbb{R}^{m}$, it holds that:
$\sum_{i,j=1}^{m}w_{i}K((x_{i},v_{i}),(x_{j},v_{j}))w_{j}\geq 0$
with strict inequality as long as all points are distinct and $(w_{i})_{i=1,\ldots,m}\neq 0$, in other words the matrix $K((x_{i},v_{i}),(x_{j},v_{j}))]_{i,j}$ is a symmetric positive definite $m\times m$ matrix. Such kernel functions can in particular be constructed, as in [17, 33], by taking the tensor product of a positive definite kernel on $\mathbb{R}^{n}$ and $S^{n-1}$, for which there are many known kernel families. If one makes the additional assumption that $K$ is a continuous function vanishing at infinity, it follows from the classical theory of kernels (c.f. [56]) that there exists a unique RKHS $W$ of functions on $\mathbb{R}^{n}\times S^{n-1}$ such that $W$ is continuously embedded into $C_{0}(\mathbb{R}^{n}\times S^{n-1})$ which is “associated” to $K$. The previous statement means specifically that the Hilbert inner product on $W$ between any two functions $f(\cdot)=K((x,v),\cdot)$ and $g(\cdot)=K((x^{\prime},v^{\prime}),\cdot)$ is explicitly given by $\langle f,g\rangle_{W}=K((x,v),(x^{\prime},v^{\prime}))$. More general continuous embeddings of $W$ into spaces $C_{0}^{s}(\mathbb{R}^{n}\times S^{n-1})$ for $s\in\mathbb{N}$ can be also recovered under additional assumptions on the regularity of the kernel $K$.
Let us now consider the regression problem (the case of classification can be treated in a similar fashion). It can be framed, in this setting, similarly to a kernel ridge regression problem:
$\min_{h\in W,\beta\in\mathbb{R}}\,\lambda\|h\|_{W}^{2}+\frac{1}{R}\sum_{k=1}^{R}\left(\langle\mu^{(k)},h\rangle+\beta-y^{(k)}\right)^{2}$ (11)
where $\lambda>0$ weighs the regularization term given by the RKHS norm $\|\cdot\|_{W}$. By direct differentiation with respect to $\beta$, one finds the linear equation satisfied by the optimal $(h_{*},\beta_{*})$:
$\beta_{*}+\frac{1}{R}\sum_{k=1}^{R}\langle\mu^{(k)},h_{*}\rangle=\bar{y}.$ (12)
where $\bar{y}=\sum_{k=1}^{R}y^{(k)}/R$ is the average value of the labels. Conversely, we may derive the explicit optimality condition on $h$. We focus on the practical situation in which each $\mu^{(k)}$ is a discrete varifold that we write $\mu^{(k)}=\sum_{i=1}^{m_{k}}w_{i}^{(k)}\delta_{(x_{i}^{(k)},v_{i}^{(k)})}$. In that case, the second term of (11) only depends on the values of $h$ at the points $(x_{i}^{(k)},v_{i}^{(k)})$. In order to remove repetitions, we shall denote by $(\tilde{x}_{i},\tilde{v}_{i})$ for $i=1,\ldots,\tilde{m}$ (with $\tilde{m}\leq m\doteq m_{1}+\ldots+m_{R}$) the subset of distinct points in $\mathbb{R}^{n}\times S^{n-1}$ among the $(x_{i}^{(k)},v_{i}^{(k)})$’s. This implies that the optimal $h^{*}$ must lie in the orthogonal subspace $W_{0}^{\bot}$ where:
$W_{0}=\{h\in W:h(\tilde{x}_{i},\tilde{v}_{i})=0\text{ for all }i=1,\ldots,\tilde{m}\}$
From standard properties of RKHS, it follows that $h_{*}$ must take the form:
$h_{*}(\cdot)=\sum_{i=1}^{\tilde{m}}\alpha_{i}K((\tilde{x}_{i},\tilde{v}_{i}),\cdot)$
where $\alpha=(\alpha_{i})\in\mathbb{R}^{\tilde{m}}$ is to be determined. Introducing $\tilde{\mathbf{K}}=[K((\tilde{x}_{i},\tilde{v}_{i}),(\tilde{x}_{j},\tilde{v}_{j}))]$ the matrix of the kernel evaluations between all pairs of points, which is positive definite, one has $\|h_{*}\|_{W}^{2}=\alpha^{T}\tilde{\mathbf{K}}\alpha$. On the other hand, we may define a second kernel matrix $\mathbf{K}=[K((x_{i}^{(k)},v_{i}^{(k)}),(\tilde{x}_{j},\tilde{v}_{j}))]\in\mathbb{R}^{m\times\tilde{m}}$ so that the energy to minimize in (11) becomes:
$J(\alpha)=\lambda\alpha^{T}\tilde{K}\alpha+\frac{1}{R}\sum_{k=1}^{R}\left((w^{(k)})^{T}(\Pi^{(k)})^{T}K\alpha+\beta-y^{(k)}\right)^{2}$
in which $w^{(k)}\in\mathbb{R}^{m_{k}}$ denotes the $k$-th block vector of $w$ and $\Pi^{(k)}\in\mathbb{R}^{m_{k}\times m}$ is the orthogonal projection that extracts the $k$-th block entries of a vector in $\mathbb{R}^{m}$. This amounts to the minimization of a quadratic function of $\alpha$ for which the optimality condition is:
$\left[\lambda\tilde{\mathbf{K}}+K^{T}\left(\frac{1}{R}\sum_{k=1}^{R}(\Pi^{(k)})^{T}w^{(k)}(w^{(k)})^{T}\Pi^{(k)}\right)K\right]\alpha_{*}+K^{T}(w/R)\beta=K^{T}(w\otimes y/R)$
where we use the notation $w\otimes y$ for the vector made of the $R$ blocks $y^{(k)}w^{(k)}$, $k=1,\ldots,R$. Ultimately, combining the above with (12), we see that the optimal $\alpha_{*},\beta_{*}$ (and thus the solution of the kernel regression problem) are obtained by solving the linear system:
$\begin{bmatrix}\lambda\tilde{\mathbf{K}}+K^{T}\left(\frac{1}{R}\sum_{k=1}^{R}(\Pi^{(k)})^{T}w^{(k)}(w^{(k)})^{T}\Pi^{(k)}\right)K&K^{T}(w/R)\\ w^{T}K/R&1\end{bmatrix}\begin{bmatrix}\alpha_{*}\\ \beta_{*}\end{bmatrix}=\begin{bmatrix}K^{T}(w\otimes y/R)\\ \bar{y}\end{bmatrix}$ (13)
We note that the above block matrix is invertible for any $\lambda>0$, as can be checked using the Schur complement criterion for block matrices. It should be pointed, however, that (13) is a very large system as its size is essentially of the order of the total number of Diracs in the whole training set which can be over 100 millions in some of the experiments of Section 4. Furthermore, the linear system may be very ill-conditioned if $\lambda$ is not properly calibrated. Thus, to be practical, this kernel based approach would need to be complemented with sparse approximation methods to reduce the dimensionality of the solution space, an example being the Nystrom scheme considered in [43].
3.2 Neural network model